OpenStudy (anonymous):

Given that the difference between the roots of the quadratic equation x^2+8x+q=0 is 4,calculate the value of q. @Miracrown

Parth (parthkohli):

Let one root be \(k\). The other one is \(k+4\).\[k + (k+4) = -8\tag{Vieta's Formula #1}\]

Parth (parthkohli):

\[k\cdot (k+4) = q\tag{Vieta's Formula #2}\]

Parth (parthkohli):

Use the first equation to figure out the value of \(k\), and the second one for \(q\).

OpenStudy (anonymous):

@ParthKohli how u get -8?

OpenStudy (er.mohd.amir):

x1+x2=-8 x1.x2=q given x1-x2=4 so x1=-2 and x2=-6 then q=-2 * -6 =12 hence q=12 is answer.

Parth (parthkohli):

\[ax^2 + bx + c = 0\]\[\text{sum of roots} = -\frac{b}{a}\]\[\text{product of roots} = \frac{c}{a}\]

OpenStudy (anonymous):

Thnx @ParthKohli and @Er.Mohd.AMIR